OpenStudy (ayonnaleflore):

What is the value of 7P3? 35 840 210 70

OpenStudy (anonymous):

i dont no but ill tag someone that i no can help @mathwizzard3

OpenStudy (ayonnaleflore):

ok thanks

OpenStudy (anonymous):

oh heres someone else who can help @rishavraj

OpenStudy (anonymous):

can u help her @rishavraj

rishavraj (rishavraj):

lol yeah :))

rishavraj (rishavraj):

\[^nP_r = \frac{ n! }{ (n - r)! }\]

OpenStudy (ayonnaleflore):

7/7-3?

rishavraj (rishavraj):

its \[\frac{ 7! }{ (7 - 3)! }\] do u know wht "!" means ???

OpenStudy (ayonnaleflore):

no clue

OpenStudy (ayonnaleflore):

so its like x1x2x3x4x5x6x7/1x2x3x4x5x6x7x1x2x3 ??

rishavraj (rishavraj):

\[7! = 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1 ~~~and~~~(7 - 3)! ~=~4!~=~4 \times 3 \times 2 \times1\]

OpenStudy (ayonnaleflore):

oh ok 5040/23=210 right

rishavraj (rishavraj):

yup :)))

OpenStudy (ayonnaleflore):

ok thanks

rishavraj (rishavraj):

u welcome :))